### v1

# A solution to Banach conjecture

Abstract.
In this paper, we first prove that the origin-symmetric star body with ellipsoidal sections is an ellipsoid. Then we give a complete proof of Banach’s isometric subspace problem in finite dimensions through the John ellipsoid.
2010 Mathematics Subject Classification:
52A20, 52A22
1. Introduction
In this paper, we establish a positive answer to the following problem for finite $n$.
Problem 1.1.
Let $(V,\|\cdot\|)$ be a normed vector space (over $\mathbb{R}$) such that for some fixed $n$, $2\leq n\leq\mbox{dim}(V)$, all $n$-dimensional linear subspaces of $V$ are isometric. Is $\|\cdot\|$ neccessarily a Eulidean norm (i.e. an inner product one)?
This problem goes back to Banach in 1932 [2, Remarks on Chapter XII] and is often referred to as the ”Banach conjecture.” The conjecture asserts that the answer is always affirmative. It has been confirmed in many, though not all, dimensions. Auerbach, Mazur, and Ulam established the conjecture for the case $n=2$ in [1]. Dvoretzky [5] later proved it for infinite-dimensional spaces $V$ and all $n\geq 2$. Gromov [7] showed that the answer is positive whenever $n$ is even or $\dim(V)\geq n+2$. Bor, Hernández-Lamoneda, Jiménez-Desantiago, and Montejano extended Gromov’s theorem to all $n$ congruent to $1$ modulo $4$, with the exception of $n=133$; see [3]. More recently, Ivanov, Mamaev, and Nordskova [8] obtained an affirmative solution for $n=3$. The problem can also be formulated for complex normed spaces; see [4], [7], [10] for further details.
Dvoretzky and Gromov reduced the problem to the situation where $\mbox{dim}(V)=n+1$ for some finite integer $n\geq 2$. Then, the problem can be restated in geometric terms as follows.
Problem 1.2.
Let $n\geq 3$ and let $K\subseteq\mathbb{R}^{n}$ be an origin-symmetric convex body. Assume that all intersections of $K$ with (n1)(n-1)-dimensional linear subspaces are linearly equivalent. Is $K$ neccessarily an ellipsoid?
The symmetry condition follows directly from Montejano in [11]: If $K\subseteq\mathbb{R}^{n}$ is a convex body and all intersections of $K$ with $n$-dimensional linear subspaces are affine equivalent, then either $K$ is symmetric with respect to 0, or $K$ is a (not necessarily centered) ellipsoid. Our principal result is the following theorem, which provides affirmative answers for every finite $n\geq 3$.
Theorem 1.1.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$, if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then $K$ is an ellipsoid.
The proof of Theorem 1.1 is divided into two parts. Assume first that $K\cap\xi_{0}$ is an (n1)(n-1)-dimensional ellipsoid. Under this assumption, $K$ itself is an ellipsoid. Thus, to prove the main theorem, we choose a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then employ the linear invariance of the John ellipsoid to derive the final result.
2. Notation and Preliminaries
This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Gardner [6] and Schneider [12].
Let $\mathbb{R}^{n}$ be the classical $n$-dimensional Euclidean space. We denote by $\|x\|$ the Euclidean norm of $x\in\mathbb{R}^{n}$. We write $B^{n}=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ for the unit ball and $S^{n-1}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ for the unit sphere. We denote $\{e_{1},e_{2},\dots,e_{n}\}$ to be the standard basis in $\mathbb{R}^{n}$. For any $\xi\in S^{n-1}$, we denote $\xi^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle=0\}$ to be a hyperplane that is perpendicular to $\xi$. Here, $\langle x,\xi\rangle$ is the usual inner product in $\mathbb{R}^{n}$. The Grassmann manifold of $k$-dimensional subspaces in $\mathbb{R}^{n}$ is denoted by $G(n,k)$. $GL(n)$ and $SL(n)$ are the general linear group and the special linear group in $\mathbb{R}^{n}$. For $\phi\in GL(n)$, we denote $|\phi|=|\det(\phi)|$ by the norm of the operator $\phi$.
A set in $\mathbb{R}^{n}$ is called convex if it contains the closed line segment joining any two of its points. A convex set is a convex body if it is compact and has a non-empty interior. A convex body is strictly convex if its boundary contains no line segments.
A compact set $L$ is called a star body if the origin $O$ is an interior point of $L$, every line through $O$ meets $L$ in a line segment, and its Minkowski gauge defined by
$\|x\|_{L}=\inf\{a\geq 0:x\in aL\}$
is a continuous function on $\mathbb{R}^{n}$.
[…]
For $H\in G(n,k)$, we call a body a revolution body with respect to $H$ if, for any special orthonormal transformation $\phi\in SO(n)$ with $\phi(H)=H$, we have $\phi(K)=K$.
Let $K\subseteq\mathbb{R}^{n}$ be a convex body. The John ellipsoid of $K$ is the ellipsoid of maximal volume contained in $K$, denoted by $J(K)$. It is clear that the John ellipsoid is invariant under linear transformations; that is, for any $\phi\in GL(n)$, $J(\phi(K))=\phi(J(K))$. When the John ellipsoid is the $n$-dimensional unit Euclidean ball $B^{n}$, the body $K$ is said to be in John position.
3. Proof of the main theorem
We begin the proof with a special case.
Theorem 3.1.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric star body. If for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ is a centered ellipsoid, then $K$ is an ellipsoid.
Lemma 3.1.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If there exists a $H\in G(n,k)$ for $1\leq k\leq n-1$ such that $K\cap H$ and $K\cap H^{\perp}$ are balls, then $K$ is an ellipsoid.
Proof.
Suppose that $K\cap H$ is a $k$-dimensional ball with a radius $r_{1}$, and $K\cap H^{\perp}$ is a (nk)(n-k)-dimensional ball with a radius $r_{2}$.
Then, for any $\eta\in\mathbb{S}^{n-1}\cap H$, $K\cap\eta^{\perp}$ is an ellipsoid where $K\cap\eta^{\perp}\cap H$ is a (k1)(k-1)-dimensional ball with a radius $r_{1}$, and $K\cap\eta^{\perp}\cap H^{\perp}$ is a (nk)(n-k)-dimensional ball with a radius $r_{2}$.
Therefore, for any $\eta\in\mathbb{S}^{n-1}\cap H$, $K\cap\eta^{\perp}$ is an ellipsoid with (k1)(k-1) axes with lengths $r_{1}$ and (nk)(n-k) axes with lengths $r_{2}$, which implies $K$ is a revolution body with respect to $H^{\perp}$; hence, $K$ is an ellipsoid. ∎
Lemma 3.2.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If we define the maximal set $M(K)$ to be
$M(K)=\{\theta\in\mathbb{S}^{n-1}:\rho_{K}(\theta)\geq\rho_{K}(\eta)\mbox{ for any }\eta\in\mathbb{S}^{n-1}\},$
then, for any linearly independent vectors $\theta_{1},\cdots,\theta_{k}\in M(K)$ with $1\leq k\leq n$, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ is a $k$-dimensional ball with a maximal radius.
Proof.
Case I. If $k=1$, $K\cap\mbox{span}(\theta_{1})$ is a $1$-dimensional ball, as $K$ is origin symmetric.
Case II. If $2\leq k\leq n-1$, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ is a $k$-dimensional ellipsoid with $k$ independent vectors $\theta_{1},\cdots,\theta_{k}$ where $\rho_{K}(\theta_{1})=\cdots=\rho_{K}(\theta_{k})$.
Hence, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ has to be a $k$-dimensional ball with a maximal radius.
Case III. If $k=n$, by Case II, for any $1\leq i\leq n$, $K\cap\,\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ is a (n1)(n-1)-dimensional ball. For any $\xi\in\mathbb{S}^{n-1}$, if $\xi^{\perp}$ coincides with $\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ for some $i$, $K\cap\xi^{\perp}$ is a (n1)(n-1)-dimensional ball with a maximal radius. If $\xi^{\perp}$ does not coincide with $\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ for all $i$, we can choose n1n-1 linearly independent vectors in $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{1}}\right)$ and $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{2}}\right)$ with $i_{1}\neq i_{2}$; otherwise, $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{1}}\right)$ and $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{2}}\right)$ coincide, which gives $i_{1}=i_{2}$.
Therefore, by Case II, $K\cap\xi^{\perp}$ has to be a $k$-dimensional ball with a maximal radius. Therefore, for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ has to be a $k$-dimensional ball with a maximal radius, which implies that $K$ is a ball. ∎
It is clear to derive the following two Corollaries from Lemma 3.2.
Corollary 3.1.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. Then the maximal set $M(k)$ is a $k$-dimensional sphere in $\mathbb{S}^{n-1}$ for some $0\leq k\leq n-1$.
Corollary 3.2.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If we define the minimal set $m(K)$ to be
$m(K)=\{\theta\in\mathbb{S}^{n-1}:\rho_{K}(\theta)\leq\rho_{K}(\eta)\mbox{ for any }\eta\in\mathbb{S}^{n-1}\},$
then, $m(k)$ is a $l$-dimensional sphere in $\mathbb{S}^{n-1}$ for some $0\leq l\leq n-1$.
Lemma 3.3.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. Then either the maximal set $M(k)$ is orthogonal to the minimal set $m(K)$ or $M(K)=m(K)=\mathbb{S}^{n-1}$.
Proof.
If $M(K)\neq m(K)$, then $M(K)\cap m(K)=\emptyset$, and thus
$\mbox{dim}(M(K))+\mbox{dim}(m(K))\leq n-2;$
otherwise, $\mbox{dim}(\mbox{span}(M(K)))+\mbox{dim}(\mbox{span}(m(K)))\geq n+1$, where $\mathbb{S}^{n-1}\cap\mbox{span}(M(K))\cap\mbox{span}(m(K))\neq\emptyset$.
Case I. If $M(K)$ is 0-dimensional and $m(K)$ is (n2)(n-2)-dimensional, we choose $\xi\in M(K)$ and consider any (n3)(n-3)-dimensional subsphere $H$ of $m(K)$.
Then $K\cap\mbox{span}(\xi,H)$ forms an ellipsoid, where the ellipsoid attains its maximum at $\xi$ and its minimum in $H$; hence $\xi\perp H$, which gives $M(K)\perp m(K)$.
Case II. If $M(K)$ is $k$-dimensional ($k\geq 1$) and $m(K)$ is $l$-dimensional ($l\leq n-3$), then for any $\xi\in M(K)$, $K\cap\,\mbox{span}(\xi,m(K))$ is an $l+2$-dimensional ellipsoid, where the ellipsoid attains its maximum at $\xi$ and its minimum in $m(K)$; hence $\xi\perp m(K)$, which gives $M(K)\perp m(K)$. ∎
Proof of Theorem 3.1.
First by Corollary 3.1 and 3.2, we can set the maximal set $M(K)$ to be $k_{1}$-dimensional and $m(K)$ to be $l_{1}$-dimensional. If $M(K)=m(K)=\mathbb{S}^{n-1}$, $K$ is a ball. If not, by Lemma 3.3, we have $M(K)\perp m(K)$ with $k_{1}+l_{1}\leq n-2$.
If $k_{1}+l_{1}=n-2$, by Lemma 3.1, we have $K$ to be ellipsoid.
If $k_{1}+l_{1}<n-2$, we can define an orthonomal bases $\{e_{1},\dots,e_{n}\}$ such that $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{1},\dots,e_{k_{1}+1}\right)=M(K)$ and $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{n-l_{1}},\dots,e_{n}\right)=m(K)$.
Now we can set $M_{1}:=\max\{\rho_{K}(\theta):\theta\in\mathbb{S}^{n-1}\}$, $m_{1}=\min\{\rho_{K}(\theta):\theta\in\mathbb{S}^{n-1}\}$, and a linear transformation $T_{1}$ to be
$\left(\begin{array}[]{ccc}\frac{m_{1}}{M_{1}}I_{k_{1}+1}&&\\ &I_{n-k_{1}-l_{1}-2}&\\ &&I_{l_{1}+1}.\end{array}\right)$
Take $K_{2}=T_{1}(K)$. It is clear that $K_{2}$ satisfies the given condition in Theorem 3.1, as $T_{1}$ is a linear transformation. Moreover,
$K_{2}\cap\mbox{span}\{e_{1},\dots,e_{k_{1}+1},e_{n-l_{1}},\dots,e_{n}\}$
is a $(k_{1}+l_{1}+2)$-dimensional ball with radius $m_{1}$.
Note that, by Corollary 3.1 and 3.2, we can set the maximal set $M(K_{2})$ to be $k_{2}$-dimensional and $m(K_{2})$ to be $l_{2}$-dimensional.
We claim $\mathbb{S}^{n-1}\cap\mbox{span}(M(K),m(K))\subseteq m(K_{2})$. Hence,
$l_{2}\geq k_{1}+l_{1}+1.$
Then, by Lemma 3.3, either $m(K_{2})=\mathbb{S}^{n-1}$ or $M(K_{2})\perp m(K_{2})$.
If $m(K_{2})=\mathbb{S}^{n-1}$, we have $K_{2}$ to be a ball which gives $K$ being an ellipsoid.
If $M(K_{2})\perp m(K_{2})$, we choose the same orthonomal bases $\{e_{1},\dots,e_{n}\}$ such that $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{k_{1}+1},\dots,e_{k_{1}+k_{2}+2}\right)=M(K_{2})$.
Now we can set $M_{2}:=\max\{\rho_{K_{2}}(\theta):\theta\in\mathbb{S}^{n-1}\}$, $m_{2}=\min\{\rho_{K_{2}}(\theta):\theta\in\mathbb{S}^{n-1}\}$ with $m_{2}=m_{1}$, and a linear transformation $T_{2}$ to be
$\left(\begin{array}[]{cccc}I_{k_{1}+1}&&&\\ &\frac{m_{1}}{M_{2}}I_{k_{2}+1}&&\\ &&I_{n-k_{1}-l_{1}-k_{2}-3}&\\ &&&I_{l_{1}+1}.\end{array}\right)$
Take $K_{3}=T_{2}(K_{2})$. It is clear that $K_{3}$ satisfies the given condition in Theorem 3.1, as $T_{2}$ is a linear transformation. Moreover,
$K_{3}\cap\mbox{span}\{e_{1},\dots,e_{k_{1}+k_{2}+2},e_{n-l_{1}},\dots,e_{n}\}$
is a $(k_{1}+k_{2}+l_{1}+3)$-dimensional ball with radius $m_{1}$.
Repeat the above process, we can finally get a ball after finite linear transformations, which means $K$ is an ellipsoid.
proof of the claim. Since $m(K_{2})\leq m_{1}$, if we assume the opposite of the claim, $m(K_{2})<m_{1}$, which means there exists $\eta\in\mathbb{S}^{n-1}$ such that $\rho_{K_{2}}(\eta)<m_{1}$ and $\eta\notin M(K)\cap m(K)$.
Thus, we choose the preimage of $\eta$ under $T_{1}$ to be $\eta$’. Since $k_{1}+l_{1}<n-2$, we have $K\cap\mbox{span}(M(K),\eta^{\prime})$ is at most $k_{1}+2$ dimensional with $k_{1}+2\leq n-1$ thus an ellipsoid with longest axes in $M(K)$ and another axis $b$. Note that $m_{1}$ is the minimum of $K$, then $b\geq m(K)$.
Now we decompose $\rho_{K}(\eta^{\prime})\eta^{\prime}$ into two parts
$\rho_{K}(\eta^{\prime})\eta^{\prime}=M_{1}\eta_{1}+\tilde{b}\eta_{2},$
where $\eta_{1}\in M(K)$ and $\eta_{2}\in\mathbb{S}^{n-1}\cap M(K)^{\perp}$.
And thus, we consider the elliptic $K\cap\mbox{span}(\eta_{1},\eta_{2})$ with long axis $M_{1}$ and short axis $b$. Then $T_{1}(K\cap\mbox{span}(\eta_{1},\eta_{2}))$ gives another elliptic with long axis $b\,(\geq m_{1})$ and short axis $m_{1}$, where $\rho_{K_{2}}(\eta)\eta\in T_{1}(K\cap\mbox{span}(\eta_{1},\eta_{2}))$ with $\rho_{K_{2}}(\eta)\geq m_{1}$, contradiction.
Now we can prove the main theorem using John ellipsoids.
Theorem 3.2.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}.$
Moreover, $E=J(K)$.
Proof.
First, we need to show the existence of the star body $E$. It is sufficient to show the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$. Since $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, for any $x\in K\cap\xi_{0}^{\perp}$, we have
$\|\phi_{\xi}(x)\|_{K}$ $=\inf\{\lambda:\phi_{\xi}(x)\in\lambda K\}$
$=\inf\{\lambda:\phi_{\xi}(x)\in\lambda(K\cap\xi^{\perp})\}$
$=\inf\{\lambda:\phi_{\xi}(x)\in\lambda(\phi_{\xi}(K\cap\xi_{0}^{\perp}))\}$
$=\inf\left\{\lambda:\phi_{\xi}(x)\in\phi_{\xi}\left(\frac{\lambda}{|\phi_{\xi}|}(K\cap\xi_{0}^{\perp})\right)\right\}$
$=\inf\left\{\lambda:x\in\frac{\lambda}{|\phi_{\xi}|}(K\cap\xi_{0}^{\perp})\right\}$
$=|\phi_{\xi}|\inf\left\{\lambda:x\in\lambda(K\cap\xi_{0}^{\perp})\right\}$
$=|\phi_{\xi}|\|x\|_{K}.$
Then the continuity of the Minkowski gauge holds
$\left|\|\phi_{\xi}(x)\|_{K}-\|\phi_{\eta}(x)\|_{K}\right|\leq\|x\|_{K}\big||\phi_{\xi}|-|\phi_{\eta}|\big|\leq c\|x\|_{K}\|\xi-\eta\|$
for some constant $c$. The second inequality follows from the continuity of the volume of the sections.
Therefore, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$ follows, which gives an origin-symmetric star body
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}.$
Then, by Theorem 3.1, $E$ is an ellipsoid in $K$ with $E\cap\xi^{\perp}$ being the John ellipsoid of $K\cap\xi^{\perp}$ for any $\xi\in\mathbb{S}^{n-1}$.
For any ellipsoid $\tilde{E}$ in $K$ and $\xi\in\mathbb{S}^{n-1}$, $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$ because $E\cap\xi^{\perp}$ is the John ellipsoid of $K\cap\xi^{\perp}$. As a result of ellipsoid being the intersection body, by Koldobsky and Zhang’s result (see [9] and [13]), we have $|\tilde{E}|\leq|E|$, which implies $E$ is the John ellipsoid of $K$. ∎
Proof of the Theorem 1.1.
By Theorem 2.2, $K$ contain a John ellipsoid $E$ where for any $\xi\in\mathbb{S}^{n-1}$, $E\cap\xi^{\perp}$ is the John ellipsoid of $K\cap\xi^{\perp}$. We can affine transform $K$ to the John’s position $\tilde{K}$, where $E$ becomes a ball $B$. Since John ellipsoid is linear invariant, $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ to be its John ellipsoid, then $\phi_{\xi}$ has to be congruent transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ is $\phi_{\xi}(B\cap\xi^{\perp})$ which is not $B\cap\xi^{\perp}$.
Then $\tilde{K}\cap\xi^{\perp}$ have some volumes for all $\xi\in\mathbb{S}^{n-1}$, and thus by Minkowski-Funk theorem, $\tilde{K}$ has to be a ball, which implies $K$ is an ellipsoid. ∎