### v2

# A solution to Banach conjecture

Abstract.
In this paper, we first prove that the origin-symmetric star body with ellipsoidal sections is an ellipsoid. Then, we provide a complete proof of Banach’s isometric subspace problem in finite dimensions through the John ellipsoid.
Key words and phrases:
Ellipsoid characterization, convex body, central sections, John ellipsoid.
2010 Mathematics Subject Classification:
52A20, 52A22
1. Introduction
In this paper, we establish a positive answer to the following problem for finite $n$.
Problem 1.1.
Let $(V,\|\cdot\|)$ be a normed vector space (over $\mathbb{R}$) such that for some fixed $n$, $2\leq n\leq\mbox{dim}(V)$, all $n$-dimensional linear subspaces of $V$ are isometric. Is $\|\cdot\|$ neccessarily a Eulidean norm (i.e. an inner product one)?
This problem goes back to Banach in 1932 [2, Remarks on Chapter XII] and is often referred to as the ”Banach conjecture.” The conjecture asserts that the answer is always affirmative. It has been confirmed in many, though not all, dimensions. Auerbach, Mazur, and Ulam established the conjecture for the case $n=2$ in [1]. Dvoretzky [5] later proved it for infinite-dimensional spaces $V$ and all $n\geq 2$. Gromov [7] showed that the answer is positive whenever $n$ is even or $\dim(V)\geq n+2$. Bor, Hernández-Lamoneda, Jiménez-Desantiago, and Montejano extended Gromov’s theorem to all $n$ congruent to $1$ modulo $4$, with the exception of $n=133$; see [3]. More recently, Ivanov, Mamaev, and Nordskova [8] obtained an affirmative solution for $n=3$. The problem can also be formulated for complex normed spaces; see [4], [7], [10] for further details.
Dvoretzky and Gromov reduced the problem to the situation where $\mbox{dim}(V)=n+1$ for some finite integer $n\geq 2$. Then, the problem can be restated in geometric terms as follows.
Problem 1.2.
Let $n\geq 3$ and let $K\subseteq\mathbb{R}^{n}$ be an origin-symmetric convex body. Assume that all intersections of $K$ with (n1)(n-1)-dimensional linear subspaces are linearly equivalent. Is $K$ neccessarily an ellipsoid?
The symmetry condition follows directly from Montejano in [11]: If $K\subseteq\mathbb{R}^{n}$ is a convex body and all intersections of $K$ with $n$-dimensional linear subspaces are affine equivalent, then either $K$ is symmetric with respect to 0, or $K$ is a (not necessarily centered) ellipsoid. Our principal result is the following theorem, which provides affirmative answers for every finite $n\geq 3$.
Theorem 1.1.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$, if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then $K$ is an ellipsoid.
The proof of Theorem 1.1 is divided into two parts. Assume first that $K\cap\xi_{0}$ is an (n1)(n-1)-dimensional ellipsoid. Under this assumption, $K$ itself is an ellipsoid. Thus, to prove the main theorem, we choose a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then employ the linear invariance of the John ellipsoid to derive the final result.
2. Notation and Preliminaries
This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Gardner [6] and Schneider [12].
Let $\mathbb{R}^{n}$ be the classical $n$-dimensional Euclidean space. We denote by $\|x\|$ the Euclidean norm of $x\in\mathbb{R}^{n}$. We write $B^{n}=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ for the unit ball and $S^{n-1}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ for the unit sphere. We denote $\{e_{1},e_{2},\dots,e_{n}\}$ to be the standard basis in $\mathbb{R}^{n}$. For any $\xi\in S^{n-1}$, we denote $\xi^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle=0\}$ to be a hyperplane that is perpendicular to $\xi$. Here, $\langle x,\xi\rangle$ is the usual inner product in $\mathbb{R}^{n}$. The Grassmann manifold of $k$-dimensional subspaces in $\mathbb{R}^{n}$ is denoted by $G(n,k)$. $GL(n)$ and $SL(n)$ are the general linear group and the special linear group in $\mathbb{R}^{n}$. For $\phi\in GL(n)$, we denote $|\phi|=|\det(\phi)|$ by the norm of the operator $\phi$.
A set in $\mathbb{R}^{n}$ is called convex if it contains the closed line segment joining any two of its points. A convex set is a convex body if it is compact and has a non-empty interior. A convex body is strictly convex if its boundary contains no line segments.
A compact set $L$ is called a star body if the origin $O$ is an interior point of $L$, every line through $O$ meets $L$ in a line segment, and its Minkowski gauge defined by
$\|x\|_{L}=\inf\{a\geq 0:x\in aL\}$
is a continuous function on $\mathbb{R}^{n}$.
[…]
For $H\in G(n,k)$, we call a body a revolution body with respect to $H$ if, for any special orthonormal transformation $\phi\in SO(n)$ with $\phi(H)=H$, we have $\phi(K)=K$.
Let $K\subseteq\mathbb{R}^{n}$ be a convex body. The John ellipsoid of $K$ is the ellipsoid of maximal volume contained in $K$, denoted by $J(K)$. It is clear that the John ellipsoid is invariant under linear transformations; that is, for any $\phi\in GL(n)$, $J(\phi(K))=\phi(J(K))$. When the John ellipsoid is the $n$-dimensional unit Euclidean ball $B^{n}$, the body $K$ is said to be in John position.
3. Proof of the main theorem
We begin the proof with a special case.
Theorem 3.1.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric star body. If for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ is a centered ellipsoid, then $K$ is an ellipsoid.
Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [13, Theorem 3.1]). The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin). Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
Now we can prove the main theorem using John ellipsoids.
Theorem 3.2.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}.$
Moreover, $E=J(K)$.
Proof.
First, we need to show the existence of the star body $E$. It is sufficient to show the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$.
Now we define a set-valued function $\lambda:\mathbb{S}^{n-1}\to GL(n)$ to be
$\lambda(\xi):=\{\phi_{\xi}\in GL(n):K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})\}.$
And we define the linear symmetry group $G_{K\cap\xi_{0}^{\perp}}$ to be
$G_{K\cap\xi_{0}^{\perp}}:=\{g\in GL(n):g(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\}$
Note that for any $\phi_{\xi},\psi_{\xi}\in\lambda(\xi)$, we have
$\psi_{\xi}^{-1}\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp},$
which gives $\phi_{\xi}\in\psi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}}$; that is
$\lambda(\xi)=\phi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}},$
for some $\phi_{\xi}\in\lambda(\xi)$. And thus, the right action of $G_{K\cap\xi_{0}^{\perp}}$ on $GL(n)$ gives the quotient $GL(n)/G_{K\cap\xi_{0}^{\perp}}$, where the set-valued function $\lambda(x)$ becomes a map $\Phi:\mathbb{S}^{n-1}\to GL(n)\backslash G_{K\cap\xi_{0}^{\perp}}$ with
$K\cap\xi^{\perp}=\Phi(\xi)(K\cap\xi_{0}^{\perp})$
Now we need to show $\Phi$ is continuous. Note that the operator norm of $\phi_{\xi}\in\Phi(\xi)$ is bounded, since
$\|\phi_{\xi}\|\leq\frac{\max_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}{\min_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}.$
Then we can find an arbitary convergent sequence $\{\xi_{n}\}_{n=1}^{\infty}$ with $\lim_{n\to\infty}\xi_{n}=\xi$ and pick $\phi_{\xi_{n}}\in\Phi(\xi_{n})$ such that there exists a subsequence $\{\xi_{n_{k}}\}_{k=1}^{\infty}$ with $\phi_{\xi_{n_{k}}}$ converging to $\phi\in GL(n)$ in operator norm. And thus, we have
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})$
$\leq$ $d_{H}(\phi(K\cap\xi_{0}),K\cap\xi_{n_{k}}^{\perp})+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
$\leq$ $d_{H}(\phi(K\cap\xi_{0}),\phi_{\xi_{n_{k}}}(K\cap\xi_{0}^{\perp}))+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
$\leq$ $c\left|\|\phi\|_{op}-\|\phi_{\xi_{n_{k}}}\|_{op}\right|+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp}).$
As $k\to\infty$, we have
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0,$
which implies $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$, that is, $\phi\in\Phi(\xi)$. Therefore, $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
This means that for any subsequence $\{\Phi(\xi_{n_{k}})\}_{k=1}^{\infty}$, there exists a sub-subsequence $\{\Phi(\xi_{n_{k_{l}}})\}_{l=1}^{\infty}$ that converges to $\Phi(\xi)$. Consequently, we conclude that $\Phi(\xi_{n})\to\Phi(\xi)$.
Moreover, for any $g\in G_{K\cap\xi_{0}^{\perp}}$, $g(J(K\cap\xi_{0}^{\perp}))=J(g(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, where the John ellipsoid of $K\cap\xi_{0}^{\perp}$ is unique. Hence, $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Therefore, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$ follows from the continuity of $\Phi(\xi)$.
Now, we need to show that for any $\xi,\eta\in\mathbb{S}^{n-1}$,
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
holds for any $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\eta^{\perp}$. It is sufficient to show for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
Here we choose $\phi_{\zeta}\in\Phi(\zeta)$ with
$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp}.$
We partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $A_{1}$ and $A_{2}$. The set $A_{1}$ consists of all $\xi$ such that, for every $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$,
$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}=g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $B$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}\neq g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$. Moreover, the closedness of $A_{1}$ follows the closedness of $G_{K\cap\xi_{0}^{\perp}}$.
For $\xi\in A$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
where
$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=J(K\cap\xi_{0}^{\perp})\cap g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
which implies
$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(g(J(K\cap\xi_{0}^{\perp}))\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
where $g(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp})\cap\phi_{\zeta}^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}))=g\circ g^{-1}(J(K\cap\xi_{0}^{\perp})\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
that is,
$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))\cap\xi_{0}^{\perp}=J(K\cap\xi_{0}^{\perp})\cap\zeta^{\perp}.$
Therefore, we obtain
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
For $\xi\in A_{2}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp},$
where
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\xi_{0}^{\perp}).$
Therefore, using the continuity of $\Phi$ and the fact that $\phi^{-1}_{\zeta}\neq g$, for any
$\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta}}{\|(1-t)\xi_{0}+t\eta_{\zeta}\|}\quad\text{with }t\in[0,1],\ \eta_{\zeta}\perp\phi^{-1}_{\zeta}(\xi_{0}^{\perp}),\ \text{and }\eta_{\zeta}\perp\xi_{0},$
there exists a linear map $\phi_{\eta}$ satisfying
$\phi_{\eta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$
and $\phi_{\eta}(\xi^{\perp})=\xi_{0}^{\perp}$. Therefore, we obtain
$\phi^{-1}_{\eta}(K\cap\xi_{0}^{\perp}\cap\eta^{\perp})=K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\eta}(\xi_{0}^{\perp}),$
and thus, $\eta\in B$. Since $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ is compact, we can choose $\zeta_{0}$ such that $\|\eta_{\zeta_{0}},\xi_{0}\|$ reaches the maximum; then $\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta_{0}}}{\|(1-t)\xi_{0}+t\eta_{\zeta_{0}}\|}$ is in $A_{2}$.
Therefore, for any open set $O\subseteq A_{2}$ and any $\xi,\eta\in O$, there exists a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$. If this were not the case, then by the continuity of $\Phi$ we could find some $\zeta\in O$ such that
$\phi_{\alpha}\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)=g\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)$
for every $\alpha\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi_{0},\zeta)$, where $g\in G_{K\cap\xi_{0}^{\perp}}$. Fixing $\xi\in O$, consider the set
$\{\eta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:K\cap\xi_{0}^{\perp}\cap\eta^{\perp}=\phi_{\eta}(K\cap\xi_{0}^{\perp}\cap\xi^{\perp}),\,\phi\in GL(n)\},$
which is closed. Consequently, for all $\xi,\eta\in\bar{O}$ there is a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$. Then $\bar{O}\subseteq A_{2}$, which shows that $A_{2}$ is closed.
By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $A_{1}$ or $A_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$. If $A_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, the origin-symmetric star body
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}$
exsits. Then, by Theorem 3.1, $E$ is an ellipsoid in $K$ with $E\cap\xi^{\perp}$ being the John ellipsoid of $K\cap\xi^{\perp}$ for any $\xi\in\mathbb{S}^{n-1}$.
If $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, then any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are mutually linearly dependent; then, using the linear dependence of all (n1)(n-1)-dimensional sections, we deduce that all (n2)(n-2)-dimensional sections of $K$ are likewise linearly dependent; therefore, by Gromov’s theorem [7], $K$ must be an ellipsoid. And hence, $E=K$ is an ellipsoid.
For any ellipsoid $\tilde{E}$ contained in $K$ and any $\xi\in\mathbb{S}^{n-1}$, we have $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, since $E\cap\xi^{\perp}$ serves as the John ellipsoid of $K\cap\xi^{\perp}$. As a result of the ellipsoid being an intersection body, Koldobsky and Zhang’s result (see [9] and [14]) yields $|\tilde{E}|\leq|E|$, and hence $E$ must be the John ellipsoid of $K$. ∎
Remark 3.1.
The proof of the continuity of $\phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ can also be found in [3, Lemma 1.5].
Proof of the Theorem 1.1.
By Theorem 3.2, the body $K$ contains a John ellipsoid $E$ such that, for every $\xi\in\mathbb{S}^{n-1}$, the section $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$. By applying an affine transformation, we can move $K$ into John’s position, denoted by $\tilde{K}$, in which $E$ is transformed into a Euclidean ball $B$. Since the John ellipsoid is invariant under linear transformation, the set $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ as its John ellipsoid. Consequently, $\phi_{\xi}$ must be a congruence transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ would be $\phi_{\xi}(B\cap\xi^{\perp})$, which differs from $B\cap\xi^{\perp}$.
Hence, $\tilde{K}\cap\xi^{\perp}$ has the same volume for every $\xi\in\mathbb{S}^{n-1}$, and therefore, by the Minkowski–Funk theorem, $\tilde{K}$ must be a ball, which in turn implies that $K$ is an ellipsoid. ∎
4. Acknowledgement
The crucial idea emerged when the author visited the Tianyuan Mathematical Research Center in Kunming.