### v3

# A solution to Banach conjecture

Abstract.
In this paper, we begin by constructing an ellipsoid derived from the John ellipsoid of the sections. We then use this framework to present a full proof of Banach’s isometric subspace problem in finite-dimensional spaces via the John ellipsoid.
Key words and phrases:
Ellipsoid characterization, convex body, central sections, John ellipsoid.
2010 Mathematics Subject Classification:
52A20, 52A22
1. Introduction
In this paper, we establish a positive answer to the following problem for finite $n$.
Problem 1.1.
Let $(V,\|\cdot\|)$ be a normed vector space (over $\mathbb{R}$) such that for some fixed $n$, $2\leq n\leq\mbox{dim}(V)$, all $n$-dimensional linear subspaces of $V$ are isometric. Is $\|\cdot\|$ neccessarily a Eulidean norm (i.e. an inner product one)?
This problem goes back to Banach in 1932 [2, Remarks on Chapter XII] and is often referred to as the ”Banach conjecture.” The conjecture asserts that the answer is always affirmative. It has been confirmed in many, though not all, dimensions. Auerbach, Mazur, and Ulam established the conjecture for the case $n=2$ in [1]. Dvoretzky [5] later proved it for infinite-dimensional spaces $V$ and all $n\geq 2$. Gromov [7] showed that the answer is positive whenever $n$ is even or $\dim(V)\geq n+2$. Bor, Hernández-Lamoneda, Jiménez-Desantiago, and Montejano extended Gromov’s theorem to all $n$ congruent to $1$ modulo $4$, with the exception of $n=133$; see [3]. More recently, Ivanov, Mamaev, and Nordskova [8] obtained an affirmative solution for $n=3$. The problem can also be formulated for complex normed spaces; see [4], [7], and [10] for further details.
Dvoretzky and Gromov reduced the problem to the situation where $\mbox{dim}(V)=n+1$ for some finite integer $n\geq 2$. Then, the problem can be restated in geometric terms as follows.
Problem 1.2.
Let $n\geq 3$ and let $K\subseteq\mathbb{R}^{n}$ be an origin-symmetric convex body. Assume that all intersections of $K$ with (n1)(n-1)-dimensional linear subspaces are linearly equivalent. Is $K$ neccessarily an ellipsoid?
The symmetry condition follows directly from Montejano in [11]: If $K\subseteq\mathbb{R}^{n}$ is a convex body and all intersections of $K$ with $n$-dimensional linear subspaces are affine equivalent, then either $K$ is symmetric with respect to 0, or $K$ is a (not necessarily centered) ellipsoid. Our principal result is the following theorem, which provides affirmative answers for every finite $n\geq 3$.
Theorem 1.1.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$, if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then $K$ is an ellipsoid.
The proof of Theorem 1.1 is divided into two steps. We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain a corresponding ellipsoid inside $K$. With this large ellipsoid at hand, we then establish the desired conclusion.
2. Notation and Preliminaries
This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Gardner [6] and Schneider [12].
Let $\mathbb{R}^{n}$ be the classical $n$-dimensional Euclidean space. We denote by $\|x\|$ the Euclidean norm of $x\in\mathbb{R}^{n}$. We write $B^{n}=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ for the unit ball and $S^{n-1}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ for the unit sphere. We denote $\{e_{1},e_{2},\dots,e_{n}\}$ to be the standard basis in $\mathbb{R}^{n}$. For any $\xi\in S^{n-1}$, we denote $\xi^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle=0\}$ to be a hyperplane that is perpendicular to $\xi$. Here, $\langle x,\xi\rangle$ is the usual inner product in $\mathbb{R}^{n}$. The Grassmann manifold of $k$-dimensional subspaces in $\mathbb{R}^{n}$ is denoted by $G(n,k)$. $GL(n)$ and $SL(n)$ are the general linear group and the special linear group in $\mathbb{R}^{n}$. For $\phi\in GL(n)$, we denote $\|\phi\|_{op}=\max_{x\in B^{n}}\|\phi(x)\|$ by the operator norm of $\phi$.
A set in $\mathbb{R}^{n}$ is called convex if it contains the closed line segment joining any two of its points. A convex set is a convex body if it is compact and has a non-empty interior. A convex body is strictly convex if its boundary contains no line segments.
A compact set $L$ is called a star body if the origin $O$ is an interior point of $L$, every line through $O$ meets $L$ in a line segment, and its Minkowski gauge defined by
$\|x\|_{L}=\inf\{a\geq 0:x\in aL\}$
is a continuous function on $\mathbb{R}^{n}$.
[…]
For $H\in G(n,k)$, we call a body a revolution body with respect to $H$ if, for any special orthonormal transformation $\phi\in SO(n)$ with $\phi(H)=H$, we have $\phi(K)=K$.
Let $K\subseteq\mathbb{R}^{n}$ be a convex body. The John ellipsoid of $K$ is the ellipsoid of maximal volume contained in $K$, denoted by $J(K)$. It is clear that the John ellipsoid is invariant under linear transformations; that is, for any $\phi\in GL(n)$, $J(\phi(K))=\phi(J(K))$. When the John ellipsoid is the $n$-dimensional unit Euclidean ball $B^{n}$, the body $K$ is said to be in John position.
3. Proof of the main theorem
We begin the proof with a special case.
Theorem 3.1.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric star body. If for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ is a centered ellipsoid, then $K$ is an ellipsoid.
Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [13, Theorem 3.1]). The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin). Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
Now we can prove the main theorem using John ellipsoids. First, we should introduce a couple of Lemmas.
Lemma 3.1.
Let $K\subseteq\mathbb{R}^{n}$ and $\phi\in GL(n)$. Then, for every $x\in\mathbb{R}^{n}$ we have
$\|x\|_{K}=\|\phi(x)\|_{\phi(K)}.$
Proof.
$\|\phi(x)\|_{\phi(K)}$ $=\inf\{\lambda:\phi(x)\in\lambda\phi(K)\}$
$=\inf\{\lambda:\phi(x)\in\phi(\lambda K)\}$
$=\inf\{\lambda:x\in\lambda K\}$
$=\|x\|_{K}.$
Lemma 3.2.
Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
$E:=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}(J(K\cap\xi_{0}^{\perp})).$
Proof.
By the definition of $E$, the set $E$ is an origin-symmetric star set whose radial function is given by
$\rho_{E}(\theta)=\max\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta):\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\}.$
Therefore, to prove that $\rho_{E}(\theta)$ is continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
Now we define a set-valued function $\lambda:\mathbb{S}^{n-1}\to GL(n)$ to be
$\lambda(\xi):=\{\phi_{\xi}\in GL(n):K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})\}.$
And we define the linear symmetry group $G_{K\cap\xi_{0}^{\perp}}$ to be
$G_{K\cap\xi_{0}^{\perp}}:=\{g\in GL(n):g(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\}$
Note that for any $\phi_{\xi},\psi_{\xi}\in\lambda(\xi)$, we have
$\psi_{\xi}^{-1}\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp},$
which gives $\phi_{\xi}\in\psi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}}$; that is
$\lambda(\xi)=\phi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}},$
for some $\phi_{\xi}\in\lambda(\xi)$. And thus, the right action of $G_{K\cap\xi_{0}^{\perp}}$ on $GL(n)$ gives the quotient $GL(n)/G_{K\cap\xi_{0}^{\perp}}$, where the set-valued function $\lambda(x)$ becomes a map $\Phi:\mathbb{S}^{n-1}\to GL(n)\backslash G_{K\cap\xi_{0}^{\perp}}$ with
$K\cap\xi^{\perp}=\Phi(\xi)(K\cap\xi_{0}^{\perp})$
Now we need to show $\Phi$ is continuous. Note that the operator norm of $\phi_{\xi}\in\Phi(\xi)$ is bounded, since
$\|\phi_{\xi}\|_{op}\leq\frac{\max_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}{\min_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}.$
We may choose an arbitrary sequence $\{\xi_{n}\}_{n=1}^{\infty}$ converging to $\xi$, i.e., with $\lim_{n\to\infty}\xi_{n}=\xi$, and for each $n$ select $\phi_{\xi_{n}}\in\Phi(\xi_{n})$. Then there exists a subsequence $\{\xi_{n_{k}}\}_{k=1}^{\infty}$ such that the corresponding operators $\phi_{\xi_{n_{k}}}$ converge in operator norm to some $\phi\in GL(n)$. And thus, we have
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})$
$\leq$ $d_{H}(\phi(K\cap\xi_{0}),K\cap\xi_{n_{k}}^{\perp})+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
$\leq$ $d_{H}(\phi(K\cap\xi_{0}),\phi_{\xi_{n_{k}}}(K\cap\xi_{0}^{\perp}))+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
$\leq$ $c\left|\|\phi\|_{op}-\|\phi_{\xi_{n_{k}}}\|_{op}\right|+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp}).$
As $k\to\infty$, we have
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0.$
This implies that $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$, and hence $\phi\in\Phi(\xi)$. Consequently, we have $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
This means that for any subsequence $\{\Phi(\xi_{n_{k}})\}_{k=1}^{\infty}$, there exists a sub-subsequence $\{\Phi(\xi_{n_{k_{l}}})\}_{l=1}^{\infty}$ that converges to $\Phi(\xi)$. Consequently, we conclude that $\Phi(\xi_{n})\to\Phi(\xi)$.
Moreover, for any $g\in G_{K\cap\xi_{0}^{\perp}}$, we have
$g(J(K\cap\xi_{0}^{\perp}))=J(g(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp}),$
since the John ellipsoid of $K\cap\xi_{0}^{\perp}$ is unique. Consequently, $J(K\cap\xi_{0}^{\perp})$ is invariant under the group $G_{K\cap\xi_{0}^{\perp}}$. Thus, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ as a function of $\xi$ follows from the continuity of $\Phi(\xi)$. ∎
Remark 3.1.
The proof of the continuity of $\phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ can also be found in [3, Lemma 1.5].
Now we consider the origin-symmetric star body $E$. The next step is to confirm that $E\cap\xi^{\perp}$ indeed recovers $\phi_{\xi}(J(E\cap\xi_{0}^{\perp}))$.
Lemma 3.3.
Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then one of the following conditions must be satisfied.
  1. (1)
    For all $\xi,\eta\in\mathbb{S}^{n-1}$ we have
    $\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
    for every $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\eta^{\perp}$.
  2. (2)
    The body $K$ is an ellipsoid.
Proof.
Condition (1) can be reformulated as follows: for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we require that
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
for every $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$, since, by choosing $\eta$ to be a basic direction, we have
$\phi_{\xi}(K\cap\xi^{\perp})=\phi_{\xi}\phi^{-1}_{\eta}(K\cap\eta^{\perp})$
for every $\xi\in\mathbb{S}^{n-1}$.
First, we introduce a function $\Gamma:\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\to\mathbb{R}_{\geq 0}$ defined by
$\Gamma(\xi)=\sup\Big\{\Big|\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|:\ \zeta,\eta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(\xi,\xi_{0})\Big\}.$
Then $\Gamma(\xi)$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Furthermore, $\Gamma(\xi)=\Gamma(\eta)$ whenever there exists a linear transformation between $K\cap\xi_{0}^{\perp}\cap\xi$ and $K\cap\xi_{0}^{\perp}\cap\eta$, as implied by Lemma 3.1.
Now fix $\xi$. For any $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$, select $\phi_{\zeta}\in\Phi(\zeta)$ such that
$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp}.$
We partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $A_{1}$ and $A_{2}$. The set $A_{1}$ consists of all $\xi$ such that, for every $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$,
$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}=g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $A_{2}$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}\neq g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$.
We once more partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $B_{1}$ and $B_{2}$. Define
$B_{1}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)=0\},\quad B_{2}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)>0\}.$
Our goal is to show that $B_{1}$ and $B_{2}$ are closed.
Note that, for $\xi\in A_{1}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
where
$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=J(K\cap\xi_{0}^{\perp})\cap g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
which implies
$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(g(J(K\cap\xi_{0}^{\perp}))\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
where $g(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp})\cap\phi_{\zeta}^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}))=g\circ g^{-1}(J(K\cap\xi_{0}^{\perp})\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
that is,
$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))\cap\xi_{0}^{\perp}=J(K\cap\xi_{0}^{\perp})\cap\zeta^{\perp}.$
Therefore, we obtain
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$. And thus, $\Gamma(\xi)=0$, so $A_{1}\subseteq B_{1}$, which implies $B_{2}\subseteq A_{2}$.
We now focus on the set $B_{2}$. For any $\xi\in B_{2}$, and therefore $\xi\in A_{2}$, and for any $\zeta,\eta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$ and $\zeta\neq\eta$, we have
$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp},$
and, in addition,
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\eta^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\eta^{\perp}).$
Hence, by exploiting the linearity of $\phi^{-1}_{\zeta}$, for any $y\in\phi^{-1}_{\zeta}(\eta^{\perp})$, we obtain $\langle\phi_{\zeta}(y),\eta\rangle=0$. Equivalently, this can be written as $y^{T}A^{T}_{\phi_{\zeta}}\eta=0$, which implies that $y\in(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$, where $A_{\phi_{\zeta}}$ denotes the matrix associated with the linear transformation $\phi_{\zeta}$. Therefore, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$. Furthermore, the set
$E_{\zeta,\xi}:=\{A^{T}_{\phi_{\zeta}}\eta:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\}$
forms an ellipse, and the set
$C_{\zeta,\xi}:=\left\{\frac{A^{T}_{\phi_{\zeta}}\eta}{|A^{T}_{\phi_{\zeta}}\eta|}:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\right\}$
gives a circle. And thus, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$ for any $\tilde{\eta}\in C_{\zeta,\xi}$. Finally, we have
[…]
for any $\tilde{\eta}\in C_{\zeta,\xi}$.
Note that for any other point $\tilde{\xi}\in B_{2}$, there is a circle $C_{\tilde{\zeta},\tilde{\xi}}$ such that
$\Gamma(\tilde{\xi})=\Gamma(\bar{\eta})$
for all $\bar{\eta}\in C_{\tilde{\zeta},\tilde{\xi}}$. Consequently, the function $\Gamma$ must be constant on $B_{2}$, since any two circles on $\mathbb{S}^{n-1}$ intersect. Therefore, $B_{2}$ is closed.
By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $B_{1}$ or $B_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$. If $B_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we have $\Gamma(\xi)=0$ for any $\xi$, and hence Condition (1) holds.
If $B_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ holds, then $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ implies that any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are all linearly dependent, because for every $\tilde{\eta}\in C_{\zeta,\xi}$ there is a linear transformation mapping $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ onto $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$. Using the fact that all (n1)(n-1)-dimensional sections are linearly dependent, we then conclude that every (n2)(n-2)-dimensional section of $K$ is also linearly dependent. Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid. ∎
To establish that $E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$, it suffices to prove that $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ and $\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))$ coincide for any $k$-dimensional subspace ($1\leq k\leq n-2$). This leads to the following lemma, which is proved by induction.
Lemma 3.4.
Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then one of the following conditions must be satisfied.
  1. (1)
    For all $H\in G(n,k)$ with $1\leq k\leq n-2$, we have
    $\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
    for every $\theta\in\mathbb{S}^{n-1}\cap H$ and $\xi,\eta\in H^{\perp}$.
  2. (2)
    The body $K$ is an ellipsoid.
Proof.
If $E$ is not an ellipsoid, then by Lemma 3.3, Condition (1) is satisfied for k=n2k=n-2. Now assume that Condition (1) holds for $k$.
For any $H\in G(n,k-1)$ and any $\xi\in H^{\perp}$, we have
$\mathrm{span}(H,\xi)\in G(n,k).$
Therefore, by the assumption of Condition (1), we obtain
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
for every $\theta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(H,\xi)$ and every $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$. In particular, this implies that
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
for all $\theta\in\mathbb{S}^{n-1}\cap H$ and all $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$, where
$\dim\bigl((\mathrm{span}(H,\xi))^{\perp}\bigr)=n-1-\dim(H)\geq 2.$
That Condition (1) is valid for k1k-1 follows from the fact that
$(\mathrm{span}(H,\xi_{1}))^{\perp}\cap(\mathrm{span}(H,\xi_{2}))^{\perp}\neq\emptyset$
for any $\xi_{1},\xi_{2}\in H^{\perp}$.
In particular, in the case $k=1$, for any $\theta\in\mathbb{S}^{n-1}$ and any $\xi,\eta\in\mathbb{S}^{n-1}\cap\theta^{\perp}$, we have
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{E}(\theta).$
Consequently,
$E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi^{\perp}).$
Now we can show the following theorem.
Theorem 3.2.
Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
$E:=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}(J(K\cap\xi_{0}^{\perp})).$
Moreover, $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$ for any $\xi\in\mathbb{S}^{n-1}$ and $E=J(K)$.
Proof.
By Lemmas 3.2 and 3.4, either $K$ itself is an ellipsoid with $E=K$, or $E$ is an origin-symmetric star body satisfying $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$. In the latter case, Theorem 3.1 implies that $E$ is an ellipsoid.
Now take an arbitrary ellipsoid $\tilde{E}$ contained in $K$. For every $\xi\in\mathbb{S}^{n-1}$, we have the inequality $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, because $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$. Since the ellipsoid is an intersection body, the theorem of Koldobsky and Zhang (see [9] and [14]) implies that $|\tilde{E}|\leq|E|$. Consequently, $E$ is indeed the John ellipsoid of $K$. ∎
Proof of the Theorem 1.1.
By Theorem 3.2, the body $K$ contains a John ellipsoid $E$ such that, for every $\xi\in\mathbb{S}^{n-1}$, the section $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$. By applying an affine transformation, we can move $K$ into John’s position, denoted by $\tilde{K}$, in which $E$ is transformed into a Euclidean ball $B$. Since the John ellipsoid is invariant under linear transformation, the set $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ as its John ellipsoid. Consequently, $\phi_{\xi}$ must be a congruence transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ would be $\phi_{\xi}(B\cap\xi^{\perp})$, which differs from $B\cap\xi^{\perp}$.
Hence, $\tilde{K}\cap\xi^{\perp}$ has the same volume for every $\xi\in\mathbb{S}^{n-1}$, and therefore, by the Minkowski–Funk theorem, $\tilde{K}$ must be a ball, which in turn implies that $K$ is an ellipsoid. ∎
4. Acknowledgement
The crucial idea emerged when the author visited the Tianyuan Mathematical Research Center in Kunming.