### v9

# A solution to Banach conjecture

Abstract.
In this paper, we begin by constructing two families of origin-symmetric star bodies derived from the John ellipsoid of the sections. We then apply the level-set method to obtain a complete proof of Banach’s isometric subspace problem in finite-dimensional spaces via the John ellipsoid.
Key words and phrases:
Ellipsoid characterization, convex body, central sections, John ellipsoid.
2010 Mathematics Subject Classification:
52A20, 52A22
1. Introduction
In this paper, we establish a positive answer to the following problem for finite $n$.
Problem 1.1.
Let $(V,\|\cdot\|)$ be a normed vector space (over $\mathbb{R}$) such that for some fixed $n$, $2\leq n\leq\mbox{dim}(V)$, all $n$-dimensional linear subspaces of $V$ are isometric. Is $\|\cdot\|$ neccessarily a Eulidean norm (i.e. an inner product one)?
This problem goes back to Banach in 1932 [2, Remarks on Chapter XII] and is often referred to as the ”Banach conjecture.” The conjecture asserts that the answer is always affirmative. It has been confirmed in many, though not all, dimensions. Auerbach, Mazur, and Ulam established the conjecture for the case $n=2$ in [1]. Dvoretzky [5] later proved it for infinite-dimensional spaces $V$ and all $n\geq 2$. Gromov [8] showed that the answer is positive whenever $n$ is even or $\dim(V)\geq n+2$. Bor, Hernández-Lamoneda, Jiménez-Desantiago, and Montejano extended Gromov’s theorem to all $n$ congruent to $1$ modulo $4$, with the exception of $n=133$; see [3]. More recently, Ivanov, Mamaev, and Nordskova [9] obtained an affirmative solution for $n=3$. This outcome leads from the analysis of sections back to the entire convex body, which provides the original insight for addressing this problem. The problem can also be formulated for complex normed spaces; see [4], [8], and [11] for further details.
Dvoretzky and Gromov reduced the problem to the situation where $\mbox{dim}(V)=n+1$ for some finite integer $n\geq 2$. Then, the problem can be restated in geometric terms as follows.
Problem 1.2.
Let $n\geq 3$ and let $K\subseteq\mathbb{R}^{n}$ be an origin-symmetric convex body. Assume that all intersections of $K$ with (n1)(n-1)-dimensional linear subspaces are linearly equivalent. Is $K$ neccessarily an ellipsoid?
The symmetry condition follows directly from Montejano in [12]: If $K\subseteq\mathbb{R}^{n}$ is a convex body and all intersections of $K$ with $n$-dimensional linear subspaces are affine equivalent, then either $K$ is symmetric with respect to 0, or $K$ is a (not necessarily centered) ellipsoid. Our principal result is the following theorem, which provides affirmative answers for every finite $n\geq 4$.
Theorem 1.1.
Let $K\subset\mathbb{R}^{n}$ with $n\geq 4$ be an origin-symmetric convex body. Fix some direction $\xi_{0}\in\mathbb{S}^{n-1}$. Suppose that for every $\xi\in\mathbb{S}^{n-1}$ there exists a linear map $\phi_{\xi}\in GL(n)$ such that
$K\cap\xi^{\perp}=\phi_{\xi}\bigl(K\cap\xi_{0}^{\perp}\bigr)$
and moreover $\phi_{\xi}(\xi_{0})=\xi$. Then $K$ must be an ellipsoid.
Indeed, Banach’s conjecture requires only a linear map on G(n,n1)G(n,n-1). However, to complete the proof, we extend this linear map to the entire space by setting
$\phi_{\xi}=\varphi_{\xi}\circ\psi_{\xi_{0},\xi},$
where $\psi_{\xi_{0},\xi}\in SO(n)$ is a rotation sending $\xi_{0}$ to $\xi$, and $\varphi_{\xi}\in GL(n)$ satisfies $\varphi_{\xi}(\xi)=\xi$. Then the proof of Theorem 1.1 is divided into three steps. We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain two families of corresponding star bodies $\{E_{\alpha}\}_{\alpha\in[0,1]}$ and $\{\tilde{E}_{\beta}\}_{\beta\in[0,1]}$ inside $K$. Secondly, we apply the level set method (see Süss’s conjecture [10, 13, 16] for further details) to demonstrate that there exists a value $t_{0}$ such that $t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}$. Having established these identities, we can then derive the desired result.
2. Notation and Preliminaries
This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Fuente [6], Gardner [7], and Schneider [14].
Let $\mathbb{R}^{n}$ denote the standard $n$-dimensional Euclidean space. For $x\in\mathbb{R}^{n}$, we write $\|x\|$ for its Euclidean norm. The unit ball is defined as $B^{n}=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$, and the unit sphere as $S^{n-1}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$. We use $\{e_{1},e_{2},\dots,e_{n}\}$ to denote the standard basis of $\mathbb{R}^{n}$. For any $\xi\in S^{n-1}$, we define
$\xi^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle=0\}$
to be the hyperplane orthogonal to $\xi$, where $\langle x,\xi\rangle$ is the usual inner product on $\mathbb{R}^{n}$. Furthermore, we set
$\xi_{+}^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle\geq 0\}\quad\text{and}\quad\xi_{-}^{\perp}:=\{x\in\mathbb{R}^{n}:\langle x,\xi\rangle\leq 0\}$
for the two half-spaces determined by the hyperplane $\xi^{\perp}$.
The Grassmann manifold of $k$-dimensional subspaces of $\mathbb{R}^{n}$ is written as $G(n,k)$. The symbols $GL(n)$ and $SL(n)$ denote, respectively, the general linear group and the special linear group on $\mathbb{R}^{n}$. For any $\phi\in GL(n)$, its operator norm is defined by $\|\phi\|_{op}=\max_{x\in B^{n}}\|\phi(x)\|$.
A subset of $\mathbb{R}^{n}$ is said to be convex if, for any two points in the set, the entire closed line segment joining them lies in the set. A convex set is called a convex body if it is compact and has non-empty interior. A convex body is strictly convex if no line segment is contained in its boundary.
A compact set $L$ is called a star body if the origin $O$ lies in the interior of $L$, each line passing through $O$ intersects $L$ in a (possibly degenerate) line segment, and its Minkowski gauge, given by
$\|x\|_{L}=\inf\{a\geq 0:x\in aL\},$
is a continuous function on $\mathbb{R}^{n}$. The radial function of $L$ is defined by $\rho_{L}(x)=\|x\|_{L}^{-1}$ for all $x\in\mathbb{R}^{n}\setminus\{O\}$. When $x$ lies on the unit sphere $S^{n-1}$, the value $\rho_{L}(x)$ represents the distance from the origin to the boundary of $L$ in the direction of $x$.
Let $K\subseteq\mathbb{R}^{n}$ be a convex body. The John ellipsoid of $K$, denoted $J(K)$, is defined as the ellipsoid of largest volume that is contained in $K$. Clearly, the John ellipsoid behaves equivariantly under linear maps: for every $\phi\in GL(n)$, we have $J(\phi(K))=\phi(J(K))$.
The level set method (see [10, 13, 16] for more details) consists in identifying a function on the sphere that remains unchanged under linear (congruent) transformations relating sections or projections, so that one of its level sets can coincide with the entire sphere.
A continuous map $f:X\to Y$ is called null-homotopic if it is homotopic to a constant map. In other words, there exists a homotopy $F(x,t):X\times[0,1]\to Y$ such that $F(x,0)=f(x)$ and $F(x,1)$ is a constant map.
On the sphere $\mathbb{S}^{n-1}$, an equator is an (n2)(n-2)-dimensional subsphere of the form $\mathbb{S}^{n-1}\cap\xi^{\perp}$ for some $\xi\in\mathbb{S}^{n-1}$. The geodesic distance between points $\theta_{1}$ and $\theta_{2}$, written $d_{\mathbb{S}^{n-1}}(\theta_{1},\theta_{2})$, is defined as the length of the shorter geodesic segment connecting them, which is equal to the angle between $\theta_{1}$ and $\theta_{2}$. The geodesic ball of radius $r$ centered at $o$ is given by
$\{\theta\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(\theta,o)\leq r\}.$
3. Proof of the main theorem
We begin the proof with a special case.
Theorem 3.1.
Let $K\subset\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric star body. If for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ is a centered ellipsoid, then $K$ is an ellipsoid.
Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [15, Theorem 3.1]). The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin). Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
Now we can prove the main theorem using John ellipsoids. First, we should introduce a couple of Lemmas.
Lemma 3.1.
Let $K\subseteq\mathbb{R}^{n}$ and $\phi\in GL(n)$. Then, for every $x\in\mathbb{R}^{n}$ we have
$\|x\|_{K}=\|\phi(x)\|_{\phi(K)}.$
Proof.
$\|\phi(x)\|_{\phi(K)}$ $=\inf\{\lambda:\phi(x)\in\lambda\phi(K)\}$
$=\inf\{\lambda:\phi(x)\in\phi(\lambda K)\}$
$=\inf\{\lambda:x\in\lambda K\}$
$=\|x\|_{K}.$
Lemma 3.2.
Let $K\subset\mathbb{R}^{n}$ with $n\geq 3$ be an origin-symmetric convex body. Fix some direction $\xi_{0}\in\mathbb{S}^{n-1}$. Suppose that for every $\xi\in\mathbb{S}^{n-1}$, there exists a linear map $\phi_{\xi}\in GL(n)$ such that
$K\cap\xi^{\perp}=\phi_{\xi}\bigl(K\cap\xi_{0}^{\perp}\bigr)$
and moreover $\phi_{\xi}(\xi_{0})=\xi$. Then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ gives rise to two families of origin-symmetric star bodies $\{E_{\alpha}\}_{\alpha\in[0,1]}$ and $\{\tilde{E}_{\beta}\}_{\beta\in[0,1]}$ defined by
$\begin{cases}E_{\alpha}:=\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cap\alpha(K\cap\xi_{0}^{\perp})\big),\\[4.0pt] \tilde{E}_{\beta}:=\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup\beta(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c}.\end{cases}$
Proof.
By the definitions of $E_{\alpha}$ and $\tilde{E}_{\beta}$, it follows that both $E_{\alpha}$ and $\tilde{E}_{\beta}$ are origin-symmetric star sets, with radial functions given by
$\rho_{E_{\alpha}}(\theta)=\sup\big\{\min\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),\alpha\rho_{K}(\theta)\}:\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\}$
and
$\rho_{\tilde{E}_{\beta}}(\theta)=\inf\big\{\max\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),\beta\rho_{K}(\theta)\}:\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\},$
respectively, here $\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi^{\perp}$. Therefore, to prove that the radial functions of $E_{\alpha}$ and $\tilde{E}_{\beta}$ are continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
Now we define a set-valued function $\lambda:\mathbb{S}^{n-1}\to GL(n)$ to be
$\lambda(\xi):=\{\phi_{\xi}\in GL(n):K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})\mbox{ and }\phi_{\xi}(\xi_{0})=\xi\}.$
And we define the linear symmetry group $G_{K\cap\xi_{0}^{\perp}}$ to be
$G_{K\cap\xi_{0}^{\perp}}:=\{g\in GL(n):g(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\mbox{ and }g(\xi_{0})=\xi_{0}\}.$
Note that for any $\phi_{\xi},\psi_{\xi}\in\lambda(\xi)$, we have
$\psi_{\xi}^{-1}\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp},$
which gives $\phi_{\xi}\in\psi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}}$; that is
$\lambda(\xi)=\phi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}},$
for some $\phi_{\xi}\in\lambda(\xi)$. And thus, the right action of $G_{K\cap\xi_{0}^{\perp}}$ on $GL(n)$ gives the quotient $GL(n)/G_{K\cap\xi_{0}^{\perp}}$, where the set-valued function $\lambda(x)$ becomes a map $\Phi:\mathbb{S}^{n-1}\to GL(n)\backslash G_{K\cap\xi_{0}^{\perp}}$ with
$K\cap\xi^{\perp}=\Phi(\xi)(K\cap\xi_{0}^{\perp})$
Now we need to show $\Phi$ is continuous. Note that the operator norm of $\phi_{\xi}\in\Phi(\xi)$ is bounded, since
$\|\phi_{\xi}\|_{op}\leq\frac{\max_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}{\min_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}.$
We may choose an arbitrary sequence $\{\xi_{n}\}_{n=1}^{\infty}$ converging to $\xi$, i.e., with $\lim_{n\to\infty}\xi_{n}=\xi$, and for each $n$ select $\phi_{\xi_{n}}\in\Phi(\xi_{n})$. Then there exists a subsequence $\{\xi_{n_{k}}\}_{k=1}^{\infty}$ such that the corresponding operators $\phi_{\xi_{n_{k}}}$ converge in operator norm to some $\phi\in GL(n)$. And thus, we have
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})$
$\leq$ $d_{H}(\phi(K\cap\xi_{0}),K\cap\xi_{n_{k}}^{\perp})+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
$\leq$ $d_{H}(\phi(K\cap\xi_{0}),\phi_{\xi_{n_{k}}}(K\cap\xi_{0}^{\perp}))+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
$\leq$ $c\left|\|\phi\|_{op}-\|\phi_{\xi_{n_{k}}}\|_{op}\right|+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp}).$
As $k\to\infty$, we have
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0.$
This implies that $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$. Moreover,
$\phi(\xi_{0})=\lim_{k\to\infty}\phi_{\xi_{n_{k}}}(\xi_{0})=\lim_{k\to\infty}\xi_{n_{k}}=\xi.$
and hence $\phi\in\Phi(\xi)$. Consequently, we have $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
This means that for any subsequence $\{\Phi(\xi_{n_{k}})\}_{k=1}^{\infty}$, there exists a sub-subsequence $\{\Phi(\xi_{n_{k_{l}}})\}_{l=1}^{\infty}$ that converges to $\Phi(\xi)$. Consequently, we conclude that $\Phi(\xi_{n})\to\Phi(\xi)$.
Moreover, for any $g\in G_{K\cap\xi_{0}^{\perp}}$, we have
$g(J(K\cap\xi_{0}^{\perp}))=J(g(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp}),$
since the John ellipsoid of $K\cap\xi_{0}^{\perp}$ is unique. Consequently, $J(K\cap\xi_{0}^{\perp})$ is invariant under the group $G_{K\cap\xi_{0}^{\perp}}$. Thus, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ as a function of $\xi$ follows from the continuity of $\Phi(\xi)$. ∎
Remark 3.1.
A proof that $\Phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ is continuous is also provided in [3, Lemma 1.5].
For a fixed $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}_{0}$, we introduce the set $\tilde{\Lambda}[\theta]\subseteq\mathbb{S}^{n-1}$ by
$\tilde{\Lambda}[\theta]:=\{\,\phi_{\xi}(\theta):\phi_{\xi}\in\Phi(\xi)\ \text{and}\ \xi\in\mathbb{S}^{n-1}\,\}.$
Observe that, for any $\xi\in\mathbb{S}^{n-1}$, Lemma 3.1 implies
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\|\eta\|_{K}$
$=$ $\big\|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\eta\big\|_{K\cap\xi^{\perp}}$
$=$ $\big\|\partial(\phi_{\xi}(J(K\cap\xi_{0}^{\perp})))\cap\{\lambda\eta:\lambda>0\}\big\|_{K\cap\xi^{\perp}}$
$=$ $\big\|\partial(J(K\cap\xi_{0}^{\perp}))\cap\{\lambda\phi^{-1}_{\xi}(\theta):\lambda>0\}\big\|_{\phi^{-1}_{\xi}(K\cap\xi^{\perp})}$
$=$ $\big\|\partial(J(K\cap\xi_{0}^{\perp}))\cap\{\lambda\phi^{-1}_{\xi}(\eta):\lambda>0\}\big\|_{K\cap\xi_{0}^{\perp}}$
$=$ $\rho_{J(K\cap\xi_{0}^{\perp})}(\phi^{-1}_{\xi}(\eta))\big\|\phi^{-1}_{\xi}(\eta)\big\|_{K}.$
And hence, for any $\eta\in\tilde{\Lambda}[\theta]$, there exists a $\xi\in\mathbb{S}^{n-1}\cap\eta^{\perp}$ such that
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\|\eta\|_{K}=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}$
Our next objective is to define a system $\Gamma$ on $\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\times[0,1]$, given by
$\Gamma:=\big\{(\eta,\xi,t)\in\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\times[0,1]:\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\|\eta\|_{K}=t\ \text{and}\ \eta\perp\xi\big\}.$
Then we can define a level set $\Lambda_{t}$ to be
$\Lambda_{t}:=\{\eta\in\mathbb{S}^{n-1}:(\eta,\xi,t)\in\Gamma\mbox{ for some }\xi\in\mathbb{S}^{n-1}\cap\eta^{\perp}\}.$ (1)
It is clear that $\tilde{\Lambda}[\theta]\subseteq\Lambda_{t}$ if $\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}=t$.
Observe that the linear symmetry group is compact ([3, Lemma 2.1]), and therefore has only finitely many connected components. We select $H\subseteq G_{K\cap\xi_{0}^{\perp}}$ to be the connected component containing the identity element. By the continuity of $\Phi$, for each $\eta\in\Lambda[\theta]$ with $\phi_{\xi}\circ g_{\xi}(\theta)=\eta$, we can choose a local continuous map $\phi_{\varphi}\circ g_{\varphi}:U_{\xi}\to GL(n)$, where $U_{\xi}$ is a neighborhood of $\xi$. Fixing $\phi_{\xi_{0}}$ to be the identity, we can construct a continuous path $\gamma$ from $\theta$ to $\eta$ by piecing together these local continuous maps. This path yields
$\phi^{-1}_{\xi}(\eta)\subseteq\{g(\theta):g\in H\}.$
We redefine
$\Lambda_{t}[\theta]:=\{$ $\phi_{\xi}(\zeta):\phi_{\xi}\in\Phi(\xi)\mbox{ is local continuous},\ \zeta\in\{g(\theta):g\in H\},$
$\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}=t,\ \text{and}\ \xi\in\mathbb{S}^{n-1}\}.$
We now present several basic propositions of $\Lambda_{t}[\theta]$ and $\Lambda_{t}$.
Proposition 3.1.
Let $K\subset\mathbb{R}^{n}$ with $n\geq 4$ be an origin-symmetric convex body. Fix some direction $\xi_{0}\in\mathbb{S}^{n-1}$. Suppose that for every $\xi\in\mathbb{S}^{n-1}$, there exists a linear map $\phi_{\xi}\in GL(n)$ such that
$K\cap\xi^{\perp}=\phi_{\xi}\bigl(K\cap\xi_{0}^{\perp}\bigr)$
and moreover $\phi_{\xi}(\xi_{0})=\xi$. Then the following propositions hold.
  1. (1)
    Either $\Lambda_{t}$ is the entire sphere, or the set of $\Lambda_{t}[\theta]$ is null-homotopic and its complement $\mathbb{S}^{n-1}\setminus\Lambda_{t}[\theta]$ consists of exactly two connected components;
  2. (2)
    Either $\Lambda_{t}$ is the entire sphere, or $\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta]$ is null-homotopic, symmetric with respect to the origin, and contains an equator;
  3. (3)
    There exists some $t_{0}\in[0,1]$ for which $\Lambda_{t_{0}}=\mathbb{S}^{n-1}$.
Proof.
First, we aim to prove that if $\Lambda_{t}$ is a proper subset of the sphere, then $\Lambda_{t}[\theta]$ is null-homotopic. In fact, if $a\notin\Lambda_{t}$, then $a\notin\Lambda_{t}$ also holds because $K$ is symmetric about the origin. Thus, $a,-a\notin\Lambda_{t}[\theta]$, and hence we can define a homotopy
$F(\xi,g,s)=\frac{s\,\phi_{\xi}\circ g(\theta)+(1-s)a}{\|s\,\phi_{\xi}\circ g(\theta)+(1-s)a\|}.$
[…]
$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o)\leq r\}.$
Observe that for any $\eta\in\Lambda_{t}[\theta]$ and any $\zeta\in\mathbb{S}^{n-1}\cap\eta^{\perp}$, the set
$\{\phi_{\vartheta}\circ g_{\vartheta}(\theta)):\vartheta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\eta,\zeta)\}\subseteq\Lambda_{t}[\theta]$
traces out a curve that passes through $\eta$ and enters both $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{+}$ and $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{-}$ (see Figure 1).
Figure 1. The set $\Lambda_{t}[\theta]$ passes through $\zeta^{\perp}$.
Consequently, every boundary point $x$ of one branch $C$ of $\mathbb{S}^{n-1}\setminus\Lambda_{t}[\theta]$ must lie on a geodesic ball that is tangent at $x$ and has radius $r_{x}<\frac{\pi}{2}$. This subsphere can be written as
$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o_{x})\leq r_{x}\}.$
Otherwise, we may prolong a curve contained in $\Lambda_{t}[\theta]$ until it intersects $C$. By the compactness of $\bar{C}$, this yields a uniform radius $r_{0}<\frac{\pi}{2}$ such that $C$ lies entirely in a geodesic ball of radius $r_{0}$. Note that $\Lambda_{t}[\theta]$ is null-homotopic and $a,-a\notin\Lambda_{t}[\theta]$. Consequently, the complement $\mathbb{S}^{n-1}\setminus\Lambda_{t}[\theta]$ can only decompose into two connected components, $C_{1}$ and $C_{2}$, which contain $a$ and a-a, respectively. Now we prove Proposition (1).
For (2), assume $\Lambda_{t}$ is not the entire sphere. Then for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we have $\Lambda_{t}[\theta]=-\Lambda_{t}[-\theta]$ by the symmetry of $K$. Proposition (1) also tells us that $\Lambda_{t}[\theta]$ and $\Lambda_{t}[-\theta]$ are null-homotopic. Consequently, we must have
$\Lambda_{t}[\theta]\cap\Lambda_{t}[-\theta]\neq\emptyset;$
for if not, then
$\Lambda_{t}[-\theta]\subseteq C_{1},$
where $C_{1}$ denotes a connected component of $\mathbb{S}^{n-1}\backslash\Lambda_{t}[\theta]$. In this situation, there exists a connected component $\tilde{C}_{1}$ of $\mathbb{S}^{n-1}\backslash\Lambda_{t}[-\theta]$ that contains $\mathbb{S}^{n-1}\backslash C_{1}$; hence $\tilde{C}_{1}$ includes a large geodesic ball of the form
$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o_{1})>r_{1}\}$
for some $r_{1}<\frac{\pi}{2}$, implying that $\tilde{C}_{1}$ contains the hemisphere $\mathbb{S}^{n-1}\cap(o_{1})^{\perp}_{-}$. This, however, contradicts the requirement that $\tilde{C}_{1}$ must lie inside a smaller geodesic ball
$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o_{1})\leq r_{2}\},$
for some $r_{2}<\frac{\pi}{2}$. Therefore, both $\Lambda_{t}[\theta]$ and $\Lambda_{t}[-\theta]$ are null-homotopic with $\Lambda_{t}[\theta]\cap\Lambda_{t}[-\theta]\neq\emptyset$ and $\Lambda_{t}[\theta]=-\Lambda_{t}[-\theta]$. And hence, $\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta]$ is null-homotopy and origin-symmetric. Therefore, $\mathbb{S}^{n-1}\setminus(\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta])$ consists of two antipodal connected components, each of which lies inside one of the geodesic balls
$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,o)\leq r\}$
and
$\{y\in\mathbb{S}^{n-1}:d_{\mathbb{S}^{n-1}}(y,-o)\leq r\},$
for some $r<\frac{\pi}{2}$, respectively. Consequently, $\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta]$ contains the equator $\mathbb{S}^{n-1}\cap o^{\perp}$. This establishes Proposition (2).
For (3), we consider the continuous function
$h(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}.$
It is evident that $h$ is Lipschitz since both $K\cap\xi_{0}^{\perp}$ and $J(K\cap\xi_{0}^{\perp})$ are convex bodies in $\xi_{0}^{\perp}$.
We now assume the opposite: that $\Lambda_{t}\neq\mathbb{S}^{n-1}$ for any $t\in[m,1]$, where
$m=\inf_{\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}}\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}.$
Note that for any connected level set of $h$, denoted by $\gamma$, if $\theta\in\gamma$, then
$\{g(\theta):g\in H\}\subseteq\gamma.$
Now we choose a connected level set of $h$, denoted by $\gamma$, that divides the sphere into finitely many connected components. If, instead, there were infinitely many connected components, then the compactness of $\mathbb{S}^{n-1}$ implies that some of them would have to be arbitrarily small, which would contradict the requirement that each component is an open set.
Thus, we may assert that there exists at least one connected component $C$ of $\mathbb{S}^{n-1}\setminus\gamma$ such that
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}\cup-\overline{C}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1}.$
Indeed, if this were not the case, then by Proposition (2), the condition $\cup\{\Lambda_{t}[\theta]:\theta\in\bar{C}\cup-\bar{C}\ \text{and}\ t\in[m,1]\}\neq\mathbb{S}^{n-1}$ would imply that
$\bigcup\{\Lambda_{t}[\theta]:\theta\in C\cup-C\ \text{and}\ t\in[m,1]\}$
is an open, null-homotopic subset of $\mathbb{S}^{n-1}$ that contains some equator, whose boundary is contained in
$\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\}$
for some $\tilde{t}\in[m,1]$. Consequently, by Proposition (2), we must have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in C\cup-C\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\};$
otherwise, the set $\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\}$ would fail to be null-homotopic. This shows that, for any connected component $C$ of $\mathbb{S}^{n-1}\setminus\gamma$, we have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\bar{C}\cup-\bar{C}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\},$
and therefore
$\mathbb{S}^{n-1}=\bigcup\{\Lambda_{t}[\theta]:\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\}\subseteq\Lambda_{\tilde{t}}$,
which contradicts our assumption.
We now consider the component $C$ for which
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}\cup-\overline{C}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1}.$
Then We introduce a partially ordered set $(\mathcal{G},\prec)$ with the following structure: for each $\alpha\in\mathcal{G}$, the curve $\gamma_{\alpha}$ is a connected level set of $\{\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:h(\theta)=t_{\alpha}\}$, and $C_{\alpha}$ denotes one of the connected components of the complement $C\setminus\gamma_{\alpha}$ whose boundary lies on $\gamma_{\alpha}$. We then declare $\alpha\prec\beta$ whenever $C_{\alpha}\subsetneq C_{\beta}$.
observe that if
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1},$
for any $\beta\succ\alpha$, we have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1}.$
On the other hand, by the previous argument, if
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}\neq\mathbb{S}^{n-1},$
we have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{t_{\beta}}[\theta]:\theta\in\gamma_{\beta}\cup-\gamma_{\beta}\},$
hence, for any $\alpha\prec\beta$,
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}\subseteq\{\Lambda_{t_{\alpha}}[\theta]:\theta\in\gamma_{\alpha}\cup-\gamma_{\alpha}\}$
$\subseteq$ $\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{t_{\beta}}[\theta]:\theta\in\gamma_{\beta}\cup-\gamma_{\beta}\}.$
We now select a totally ordered subset of $\mathcal{G}$, which we denote by $\mathcal{H}$. Define
$\mathcal{F}:=\Bigl\{\alpha\in\mathcal{H}:\ \bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha},\ t\in[m,1]\}=\mathbb{S}^{n-1}\Bigr\}.$
We aim to show this set coincides with $\mathcal{H}$, i.e. $\mathcal{F}=\mathcal{H}$. Otherwise, there exists some $\beta_{0}\in\mathcal{H}$ such that for every $\alpha\prec\beta_{0}$ we have $\alpha\notin\mathcal{F}$, while for every $\beta\succ\beta_{0}$ we have $\beta\in\mathcal{F}$. However, for every $\alpha\prec\beta_{0}$ we have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{t_{\alpha}}[\theta]:\theta\in\gamma_{\alpha}\cup-\gamma_{\alpha}\},$
and therefore
$\bigcup\{\Lambda_{t}[\theta]:\theta\in C_{\beta_{0}}\cup-C_{\beta_{0}}\ \text{and}\ t\in[m,1]\}$
$\subseteq$ $\bigcup_{\alpha\prec\beta_{0}}\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}$
$\subseteq$ $\bigcup_{\alpha\prec\beta_{0}}\{\Lambda_{t_{\alpha}}[\theta]:\theta\in\gamma_{\alpha}\cup-\gamma_{\alpha}\}$
$\subseteq$ $\bigcup\{\Lambda_{t_{\beta_{0}}}[\theta]:\theta\in\gamma_{\beta_{0}}\cup-\gamma_{\beta_{0}}\},$
which yields
$\mathbb{S}^{n-1}$ $=\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta_{0}}\cup-\overline{C}_{\beta_{0}}\ \text{and}\ t\in[m,1]\}$
$\subseteq\bigcup\{\Lambda_{t_{\beta_{0}}}[\theta]:\theta\in\gamma_{\beta_{0}}\cup-\gamma_{\beta_{0}}\}\subseteq\Lambda_{\beta_{0}},$
contradicting the assumption.
Now define $\alpha_{0}$ as the infimum of $(\mathcal{H},\prec)$. Clearly, $C_{\alpha_{0}}=\emptyset$, and $\gamma_{\alpha_{0}}$ then satisfies
$\mathbb{S}^{n-1}$ $=\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha_{0}}\cup-\overline{C}_{\alpha_{0}}\ \text{and}\ t\in[m,1]\}$
$=\bigcup\{\Lambda_{t_{\alpha_{0}}}[\theta]:\theta\in\gamma_{\alpha_{0}}\cup-\gamma_{\alpha_{0}}\}\subseteq\Lambda_{\alpha_{0}},$
which contradicts our initial assumption.
Consequently, our assumption does not hold, and therefore Proposition (3) is valid. ∎
Lemma 3.3.
Let $K\subset\mathbb{R}^{n}$ with $n\geq 4$ be an origin-symmetric convex body. Fix some direction $\xi_{0}\in\mathbb{S}^{n-1}$. Suppose that for every $\xi\in\mathbb{S}^{n-1}$, there exists a linear map $\phi_{\xi}\in GL(n)$ such that
$K\cap\xi^{\perp}=\phi_{\xi}\bigl(K\cap\xi_{0}^{\perp}\bigr)$
and moreover $\phi_{\xi}(\xi_{0})=\xi$. Then there exists $t_{0}\in[0,1]$ such that
$t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}.$
Proof.
By Proposition 3.1 (3), there exists some $t_{0}\in[0,1]$ such that $\Lambda_{t_{0}}=\mathbb{S}^{n-1}$. Hence, for every $\theta\in\mathbb{S}^{n-1}$, there is a $\zeta\in\mathbb{S}^{n-1}\cap\theta^{\perp}$ satisfying
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}=t_{0}.$
Consequently,
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=t_{0}\rho_{K}(\theta).$
Using the definitions of $E_{t_{0}}$ and $\tilde{E}_{t_{0}}$, we obtain
$\rho_{E_{t_{0}}}(\theta)$ $=\sup\big\{\min\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),t_{0}\rho_{K}(\theta)\}:\ \xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\}$
$=\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=t_{0}\rho_{K}(\theta).$
Similarly,
$\rho_{\tilde{E}_{t_{0}}}(\theta)$ $=\inf\big\{\max\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),t_{0}\rho_{K}(\theta)\}:\ \xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\}$
$=\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=t_{0}\rho_{K}(\theta).$
Thus, we conclude that
$t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}.$
Now we can prove the main theorem.
Proof of the Theorem 1.1.
By Lemma 3.3, there exists $t_{0}\in[0,1]$ such that
$t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}.$
On the one hand, we observe that
$E_{t}$ $=t_{0}K$
$=\bigcup_{\xi\in\mathbb{S}^{n-1}}t_{0}\phi_{\xi}\big(K\cap\xi_{0}^{\perp}\big)$
$=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(t_{0}(K\cap\xi_{0}^{\perp})\big)$
[…]
from which it follows that
[…]
On the other hand, we obtain
$t_{0}K=$ $\tilde{E}_{t}$
$=$ $\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup t_{0}(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c}$
[…]
$=$ $\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(t_{0}(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c}$
[…]
$=$ $t_{0}K\bigcup\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\backslash t_{0}(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c},$
which implies that
$J(K\cap\xi_{0}^{\perp})\backslash t_{0}(K\cap\xi_{0}^{\perp})=\emptyset.$
Consequently,
$J(K\cap\xi_{0}^{\perp})=t_{0}(K\cap\xi_{0}^{\perp}),$
so $K\cap\xi_{0}^{\perp}$ is an ellipsoid, and by Theorem 3.1 this implies that $K$ itself is an ellipsoid. ∎